OpenStudy (bbbbbbb):

what is the greatest common factor of 10a^3b,18ab^2, and 20a?

OpenStudy (anonymous):

2a

OpenStudy (raden):

10 = 2 * 5 18 = 2 * 3^2 20 = 2^2 * 5 the GCM of them is 2 while the GCM of a^3b, ab^2, and a is only a join up both, it is 2a